OpenStudy (anonymous):

Find the exact value of the expression in simplest form: the square root of 3 times x + 3 the square root of 3 times x − 2 the square root of 3 times x − 10 I tried this, but I think I messed part of it up.

OpenStudy (anonymous):

@phi

OpenStudy (phi):

can you use the equation editor on this ? It 's too hard to decipher

OpenStudy (anonymous):

Yes, one second.

OpenStudy (anonymous):

\[\sqrt{3x}+3\sqrt{3x}-2\sqrt{3x}-10\]

OpenStudy (phi):

focus just on the terms with the sqrt(3x) can you combine them ?

OpenStudy (phi):

one ugly plus 3 more uglies take away 2 uglies. How many uglies (how many \( \sqrt{3x} \)?)

OpenStudy (anonymous):

sqrt9x ?

OpenStudy (phi):

no, all you do is count up uglies (count up the number of \(\sqrt{3x}\) ) you don't change the \(\sqrt{3x}\) only the number of them

OpenStudy (phi):

if you think of the \(\sqrt{3x}\) as cows, it would be 1 cow + 3 cows - 2 cows how many cows ? you get the same number of \(\sqrt{3x}\)

OpenStudy (anonymous):

2 sqrt 3x 's

OpenStudy (phi):

yes, you started with \[ \sqrt{3x}+3\sqrt{3x}-2\sqrt{3x}-10 \] and now have \[ 2\sqrt{3x}-10 \] which may be the answer. Or you can factor a 2 from each term to get \[2(\sqrt{3x}-5 )\]

OpenStudy (anonymous):

Awesome, thanks ^_^ I have one more, if you don't mind. I'll post it in a second.